OpenStudy (anonymous):

10x-8y=-2 9x+8y=59 A. (7, 8) B. (3, 4) C. (3, 3) D. (7, 9)

OpenStudy (anonymous):

You again? Using same sum rule we sum up these 2 equations and eliminate y 19x = 57 x = 57/19 = 3 9 * 3 + 8y = 59 27 + 8y = 59 8y = 59-27 = 32 y = 32/8 = 4 Check: 10*3 - 8 * 4 = -2 30 -32 = -2 -2 = -2 Variant B (3 4)

OpenStudy (anonymous):

You are having a math exam?